In January of year 1, Omega co bought a machine for $ 50,000. Transport costs for the machine were $ 5,000 and with the terms FOB destination and with the purchase of the machine, Omega co operating costs increased by $ 2,500 a year. The gross value of the machine is $ 10,000. but its lifespan is 5 years.
What were the depreciation of year 3 compared to double linear depreciation?
a. $ 5760
b. $ 7920
c. $ 7560
d. $ 7200
Answer #1

Answer is d. $7,200

Calculations:

Given that FOB destination. It means Freight has to be paid by seller. Hence it is not included in the cost of machine. Therefore the cost of the machine is $50,000 only.

Straight line rate = (cost ÷ useful life)/cost = ($50,000/5 years)/$50,000 = 10,000/50,000 = 0.20 or 20%

Depreciation rate under double linear = Straight line rate x 2 = 20% x 2 = 40%

Year Beginning
Book value
Depreciation
Rate
Depreciation
Expense
Ending
Book value
1 50000 x 40% = 20,000 30,000
2 30,000 x 40% = 12,000 18,000
3 18,000 x 40% = 7,200 10,800

Ending book value = Beginning book value - Depreciation expense

Beginning book value = Previous year's ending book value